# A Formal Comparison Between Chain-of-Thought and Latent Thought

### v1

@@ 3.2 CoT Suffices with Size-scaled Steps and Looped TF Suffices with Depth-scaled Loops
Assumption 3.4 (Polynomially-efficient approximation, cf. (Feng et al., 2023)).
Informally, each node function of $G_{n}$ can be approximated by a log-precision feedforward network whose parameter size is polynomial in the input length and the approximation error. We denote by $\mathrm{ff\_param}(G_{n})$ an upper bound such that every $f\in\mathcal{F}$ admits such a network with at most $\mathrm{ff\_param}(G_{n})$ parameters.
We first consider CoT. Under Assumptions 3.3, 3.4, a CoT model can simulate the computation by sequentially decoding node by node. Intermediate tokens act as scratchpad memory to store partial results, allowing the model to evaluate any node once all its predecessors are available.
@@
Let $\{G_{n}\}_{n\in\mathbb{N}}$ be a family of computation graphs that satisfy Assumptions 3.3 and 3.4. Then, for each $n\in\mathbb{N}$, there exists a log-precision CoT with parameter size bounded by $O(\mathrm{ff\_param}(G_{n}))$, such that for every input $x\in\Sigma^{n}$, the model outputs $F_{G_{n}}(x)$ in a number of steps proportional to the “size” of the graph, $\mathrm{size}(G_{n})$.
@@
The construction generalizes mechanisms in (Feng et al., 2023; Li et al., 2024): a CoT model can evaluate a graph by processing its nodes in topological order. At each step, the attention mechanism retrieves the outputs of the predecessor nodes, which are available as previously generated tokens. A feed-forward network is then applied to compute the output of the current node, and this value is generated as the next token. The iterative process yields the final result in $O(|V_{n}|)$ steps. ∎
Unlike CoT, Looped TF can simulate the computation in parallel, layer by layer, where all nodes at the same depth are computed simultaneously, provided the embedding size is sufficiently large.
Assumption 3.6 (Output bound).
The computation graph contains at most $n$ output nodes.
Theorem 3.7 (Looped TF for DAGs).
Let $\{G_{n}\}_{n\in\mathbb{N}}$ be a family of computation graphs that satisfy Assumptions 3.33.4, and 3.6. Then, for each $n\in\mathbb{N}$, there exists a log-precision Looped TF with parameter size $O(\mathrm{ff\_param}(G_{n})\cdot\mathrm{size}(G_{n}))$, such that for every input $x\in\Sigma^{n}$, it computes $F_{G_{n}}(x)$ in a number of loops proportional to the “depth” of the graph $G_{n}$, , $\mathrm{depth}(G_{n})$.
@@
The role assignment of each layer is based on Li et al. (2024), as illustrated in Figure 5. A uniform attention layer aggregates all inputs from the entire sequence into a single position. The hidden state is partitioned into dedicated blocks, each block corresponding to a node in the computation graph, and the looped feedforward layer simultaneously computes the local function of every node on its respective block. By iterating this process, all nodes at the same depth of the graph are evaluated in parallel, and successive iterations propagate the computation layer by layer. ∎
@@
Formal proofs are given in Appendix C, with illustrations in Fig. 1. These results show complementary strengths: CoT can employ intermediate steps as scratchpad memory, supporting general computation without constraints. In contrast, Looped TF can exploit structural parallelism to achieve greater efficiency, but requires additional resources.
3.3 Separation in Polylogarithmic Time
We now analyze the complexity classes of the models to establish the separation. At the beginning, we define the classes $\mathsf{CoT}$ and $\mathsf{Loop}$, following the definition of $\mathsf{CoT}$ in Li et al. (2024).
Definition 3.8 ($\mathsf{CoT}$ and $\mathsf{Loop}$).
Let $\mathsf{CoT}[T(n),d(n),s(n)]$ (resp. $\mathsf{Loop}[T(n),d(n),s(n)]$) denote the set of languages ${\mathcal{L}}:\{0,1\}^{*}\to\{0,1\}$ for which there exists a deterministic CoT (resp. Looped TF) $M_{n}$ for each input size $n$, with embedding size $O(d(n))$ and $O(s(n))$ bits of precision, such that for all $x\in\{0,1\}^{n}$, the final output token, after $O(T(n))$ decode steps (resp. loops), is ${\mathcal{L}}(x)$.
To connect these models with classical complexity-theoretic notions, we employ Boolean circuits.
Definition 3.9 (Informal: Boolean Circuits).
A Boolean circuit is a DAG over the alphabet $\Sigma=\{0,1\}$, where each internal node (gate) computes a Boolean function such as AND, OR, or NOT.
While CoT has been characterized in terms of circuit size, we show that Looped TF corresponds to the depth. Here, $\mathsf{SIZE}[s(n)]$ (resp. $\mathsf{DEPTH}[d(n)]$) denotes the class of languages decidable by (non-uniform) circuit families $\{C_{n}\}$ with size at most $O(s(n))$ (resp. depth at most $O(d(n))$).
@@ 3.3 Separation in Polylogarithmic Time
$\forall T(n)\in\mathrm{poly}(n),\,\mathsf{SIZE}[T(n)]\subseteq\mathsf{CoT}[T(n),\log{n},1].$
@@
$\forall T(n)\in\mathrm{poly}(n),\,\mathsf{DEPTH}[T(n)]\subseteq\mathsf{Loop}[T(n),\mathrm{size}(C_{n}),1].$
Since a Boolean circuit is a DAG, the claim follows from Theorem 3.7. To establish a separation, we next focus on classes of problems efficiently solvable in parallel computation.
Definition 3.12 ($\mathsf{NC}^{k}$, $\mathsf{AC}^{k}$, and $\mathsf{TC}^{k}$).
For each $k\in\mathbb{N}$, the classes $\mathsf{NC}^{k}$, $\mathsf{AC}^{k}$, and $\mathsf{TC}^{k}$ consist of languages decidable by non-uniform circuit families of size $\mathsf{poly}(n)$ and depth $O(\log^{k}n)$, using bounded-fanin Boolean gates, unbounded-fanin $\mathsf{AND}$$\mathsf{OR}$ gates, and threshold gates, respectively. It holds that $\mathsf{NC}^{k}\subseteq\mathsf{AC}^{k}\subseteq\mathsf{TC}^{k}\subseteq\mathsf{NC}^{k+1}$ (Allender et al., 1999).
Parallel computation corresponds to polylogarithmic time using a polynomial number of processors (gates) (Stockmeyer & Vishkin, 1984). Thus, we also assume polylogarithmic steps (for CoT) or loops (for Looped TF) and a polynomial embedding size. We now provide the exact power of Looped TFs in this regime.
@@
$\forall k\in\mathbb{N},\,\mathsf{Loop}[\log^{k}n,\,\mathsf{poly}(n),\,1\ \textup{(resp.\ }\log n)]=\mathsf{AC}^{k}\ \textup{(resp.\ }\mathsf{TC}^{k}).$
@@
The inclusion from circuits to Looped TFs follows from Theorem 3.7. For the converse inclusion, we build on the arguments of Li et al. (2024), which showed that under finite precision, $\mathsf{CoT}[0,\,\mathsf{poly}(n),\,1\ \textup{(resp.\ }\log n)]\subseteq\mathsf{AC}^{k}\ \textup{(resp.\ }\mathsf{TC}^{k})$. We extended this to the looped setting, yielding the desired inclusion. The formal proof is provided in Section C.6. ∎
Conversely, we establish an upper bound on CoT with a polylogarithmic number of steps.
@@
$\forall k\in\mathbb{N},\,\mathsf{CoT}[\log^{k}{n},\mathsf{poly}(n),\log{n}]\subseteq\mathsf{TC}^{k-1}.$
@@
The total $\log^{k}n$ steps can be divided into $\log^{k-1}n$ blocks, each consisting of $\log n$ steps. Since $\mathsf{CoT}[\log n,\mathsf{poly}(n),\log n]\subseteq\mathsf{TC}^{0}$ (Li et al., 2024), each block with the previous block’s outputs fed as inputs to the next block can be simulated in $\mathsf{TC}^{0}$; iterating this over $\log^{k-1}n$ layers yields a depth $O(\log^{k-1}n)$ threshold circuit, i.e., in $\mathsf{TC}^{k-1}$. ∎
This leads to the separation in their expressive power under standard complexity assumptions.
@@
$\mathsf{TC}^{k-1}\subsetneq\mathsf{NC}^{k}\Rightarrow\mathsf{CoT}[\log^{k}{n},\mathsf{poly}(n),\log{n}]\subsetneq\mathsf{Loop}[\log^{k}{n},\mathsf{poly}(n),1].$
@@
$\mathsf{TC}^{k-1}\subsetneq\mathsf{TC}^{k}\Rightarrow\mathsf{CoT}[\log^{k}{n},\mathsf{poly}(n),\log{n}]\subsetneq\mathsf{Loop}[\log^{k}{n},\mathsf{poly}(n),\log{n}].$
@@
The claims follow directly from Theorem 3.13 and Lemma 3.14. These results show that, under the same polylogarithmic number of steps or loops, Looped TF can simulate parallel computations more efficiently than CoT. This highlights the inherent parallelism of looped models.
4 Chain-of-Thought Enables Randomized Approximation
While the previous section established the efficiency of Looped TF, this section reveals the converse situation: CoT is indispensable for approximating computationally hard problems. Section 4.1 presents the preliminaries, Section 4.2 formalizes the problem setting of approximate sampling, and Section 4.3 proves a conditional separation result highlighting the superiority of CoT.
4.1 Preliminaries: Randomized Approximation Scheme for Hard Problems
While the previous section focused on problems solvable efficiently, some important problems are computationally hard and intractable in polynomial time unless $\mathsf{P}=\mathsf{NP}$. When exact computation of a function $f:\Sigma^{*}\to\mathbb{N}$ is intractable, approximation schemes are employed (Hochba, 1997).
Definition 4.1 (FPTAS).
An algorithm is a fully polynomial-time approximation scheme (FPTAS) for $f$ if, for any input $x$ and error parameter $\varepsilon>0$, it outputs an estimate $\hat{f}(x)$ satisfying $(1-\varepsilon)f(x)\leq\hat{f}(x)\leq(1+\varepsilon)f(x),$ and runs in time polynomial in $|x|$ and $1/\varepsilon$.
However, for some problems, even such an approximation remains intractable. An important development was the introduction of randomized algorithms (Karp & Luby, 1983), which enable tractable solutions by allowing a small probability of error specified by a confidence parameter.
@@ 4.1 Preliminaries: Randomized Approximation Scheme for Hard Problems
An algorithm is a fully polynomial-time randomized approximation scheme (FPRAS) for $f$ if, for any $\varepsilon>0$ and $\delta>0$, it outputs an estimate $\hat{f}(x)$ $\textit{s}.\textit{t}.$
@@
For example, the knapsack problem admits an FPTAS using a dynamic programming algorithm (Vazirani, 2001), and the DNF counting problem admits an FPRAS, where randomization through Monte Carlo methods is essential to achieve an efficient approximation (Karp et al., 1989).
Assumptions on complexity.
Although it has been conjectured that randomized computation offers no advantage over deterministic computation, i.e., $\mathsf{P}=\mathsf{BPP}$ (Nisan & Wigderson, 1994), randomization can nonetheless reduce the complexity of certain approximation algorithms (Stockmeyer, 1983). Moreover, there exist canonical problems for which only an FPRAS is currently known (Jerrum et al., 2004). We therefore present our results under the assumption that $\mathsf{FPTAS}\subsetneq\mathsf{FPRAS}$, where each denotes the set of problems that admit an FPTAS and an FPRAS, respectively.
4.2 Problem Setting
We study the problem of sampling, with a particular focus on the uniform generation of combinatorial structures. Formally, let $R\subseteq\Sigma^{*}\times\Sigma^{*}$ be a relation. For a problem instance $x\in\Sigma^{*}$, the corresponding solution set is $R(x)=\{\,y\in\Sigma^{*}:(x,y)\in R\,\}.$ The task of uniform generation is then to sample $y$ uniformly from $R(x)$. Building on prior work on the exact representation of probability distributions over strings (Nowak et al., 2024), we consider a distribution whose combinatorial structure makes exact sampling intractable, motivating the study of approximate sampling.
Definition 4.3 (FPAUS).
A fully polynomial almost uniform sampler (FPAUS) is a randomized algorithm that, given any input $x\in\Sigma^{*}$ and error tolerance $\varepsilon>0$, runs in time polynomial in $|x|$ and $\log(1/\varepsilon)$, and outputs a sample from a distribution $q(\cdot|x)$ supported on $\Sigma^{*}$ such that
where $U(R(x))$ denotes the uniform distribution over $R(x)$, extended to $\Sigma^{*}$ by assigning zero probability mass outside $R(x)$, and $\|\cdot\|_{\mathrm{TV}}$ denotes the total variation distance.
We focus on problems with a compositional structure, formalized as self-reducibility (Schnorr, 1976), where solutions can be constructed recursively from smaller subproblems.
Definition 4.4 (Informal: Self-reducibility (Schnorr, 1976)).
A relation $R$ is self-reducible if there exists a polynomial-time procedure that, given any input $x$ and prefix $y_{1:k}$ (with respect to a fixed output order), produces a sub-instance $\psi(x,y_{1:k})$ such that every solution $z$ of $\psi(x,y_{1:k})$ extends $y_{1:k}$ to a solution of $R(x)$ (and conversely), i.e., $\mathrm{solution}\bigl(\psi(x,y_{1:k})\bigr)=\{z\mid\ \mathrm{concat}(y_{1:k},z)\in R(x)\,\}.$
For computational models, we define CoT and Looped TF in the setting of random generation of strings $(y_{1},\ldots,y_{m})\in\Sigma^{m}$ via next-token prediction. Specifically, CoT stochastically generates intermediate reasoning steps before producing the target symbol $y_{i}$. In contrast, Looped TF performs additional iterations in the latent space and samples the target symbol $y_{i}$ after the final loop iteration, as in practice (Csordás et al., 2024; Snell et al., 2025; Bae et al., 2025). Definitions are in Section D.1.

### v2

@@ 3.2 CoT Suffices with Size-scaled Steps and Latent Thought Suffices with Depth-scaled Iterations
Assumption 3.4 (Poly-efficient approximation, cf. (Feng et al., 2023)).
Each node function of $G_{n}$ can be approximated by a log-precision feedforward network whose parameter size is polynomial in the input length and the inverse of the approximation error. We denote by $\mathrm{ff\_param}(G_{n})$ an upper bound such that every $f\in\mathcal{F}$ admits such a network with at most $\mathrm{ff\_param}(G_{n})$ parameters.
Under these assumptions, we show that CoT can simulate computation by sequentially decoding nodes, where intermediate tokens serve as a scratchpad allowing the evaluation of each node once all its predecessors have been computed.
@@
Let $\{G_{n}\}_{n\in\mathbb{N}}$ be a family of computation graphs that satisfy Assumptions 3.3 and 3.4. Then, for each $n\in\mathbb{N}$, there exists a log-precision CoT with parameter size bounded by $O(\mathrm{ff\_param}(G_{n}))$, such that for every input $x\in\Sigma^{n}$, the model outputs $F_{G_{n}}(x)$ with steps proportional to the “size” of the graph, i.e., $O(\mathrm{size}(G_{n}))$.
@@
At each step, the attention layer retrieves the outputs of predecessor nodes from previously generated tokens, and a feed-forward layer then computes the node function, whose output is generated as the next token. ∎
In contrast, latent thought can operate in parallel, layer by layer, where all nodes at the same depth are computed simultaneously, provided that the model has sufficient size.
Theorem 3.6 (Latent thought for DAGs).
Let $\{G_{n}\}_{n\in\mathbb{N}}$ be a family of computation graphs that satisfy Assumptions 3.3 and 3.4. Then, for each $n\in\mathbb{N}$, there exists a log-precision Coconut and looped TF with parameter size $O(\mathrm{ff\_param}(G_{n})\cdot\mathrm{size}(G_{n}))$, such that for every input $x\in\Sigma^{n}$, it computes $F_{G_{n}}(x)$ with iterations proportional to the “depth” of the graph $G_{n}$, i.e., $O(\mathrm{depth}(G_{n}))$.
@@
The role assignment of each layer is based on (Li et al., 2024), as shown in Figure 9. An attention layer aggregates its inputs into a single hidden state. Unlike discrete tokens, continuous latent states allow the simultaneous encoding of the outputs of multiple nodes, enabling the feed-forward layer to compute node functions in parallel. ∎
@@
Illustrations are provided in Fig. 2, with formal proofs deferred to Appendix B. These results reveal distinct characteristics: CoT can utilize intermediate steps as scratchpad memory and perform computations sequentially, whereas latent thought can leverage structural parallelism to achieve greater efficiency with sufficient resources.
3.3 Separation in Polylogarithmic Iterations
In this section, we shift to formal decision problems to precisely characterize the computational power of each reasoning paradigm, clarify what cannot be achieved, and use these limitations to derive rigorous separations. We begin by defining their complexity classes, as in (Li et al., 2024).
Definition 3.7 (Complexity Classes $\mathsf{CoT}$, $\mathsf{CT}$ and $\mathsf{Loop}$).
Let $\mathsf{CoT}[T(n),d(n),s(n)]$, $\mathsf{CT}[T(n),d(n),s(n)]$, and $\mathsf{Loop}[T(n),d(n),s(n)]$ denote the sets of languages ${\mathcal{L}}:\{0,1\}^{*}\to\{0,1\}$ for which there exists a deterministic CoT, Coconut, and looped TF, respectively, denoted by $M_{n}$ for each input size $n$, with embedding size $O(d(n))$ and $O(s(n))$ bits of precision, such that for all $x\in\{0,1\}^{n}$, the final output token after $O(T(n))$ iterations equals ${\mathcal{L}}(x)$.
Boolean circuits serve as a standard formal model of computation, where processes are defined by the evaluation of DAGs with well-established complexity measures.
Definition 3.8 (Informal).
A Boolean circuit is a DAG over the alphabet $\Sigma=\{0,1\}$, where each internal node (gate) computes a Boolean function such as AND, OR, or NOT. $\mathsf{SIZE}[s(n)]$ and $\mathsf{DEPTH}[d(n)]$ denote the class of languages decidable by a non-uniform circuit family $\{C_{n}\}$ with size $O(s(n))$ and depth $O(d(n))$, respectively.
First, we formalize the results of the previous section to show that latent thought iterations can represent circuit depth, whereas CoT corresponds to circuit size.
@@ 3.3 Separation in Polylogarithmic Iterations
$\forall T(n)\in\mathrm{poly}(n),$
$\mathsf{SIZE}[T(n)]\subseteq\mathsf{CoT}[T(n),\log{n},1].$
@@
For any function $T(n)\in\mathrm{poly}(n)$ and any non-uniform circuit family $\{C_{n}\}$, it holds that
$\mathsf{DEPTH}[T(n)]$ $\subseteq\mathsf{Loop}[T(n),\mathrm{size}(C_{n}),1],$
$\mathsf{DEPTH}[T(n)]$ $\subseteq\mathsf{CT}[T(n),\mathrm{size}(C_{n}),1].$
Boolean circuits serve as a formal model of parallel computations that run in polylogarithmic time using a polynomial number of processors (Stockmeyer and Vishkin, 1984).
Definition 3.11.
For each $k\in\mathbb{N}$, the classes $\mathsf{NC}^{k}$, $\mathsf{AC}^{k}$, and $\mathsf{TC}^{k}$ consist of languages decidable by non-uniform circuit families of size $\mathsf{poly}(n)$ and depth $O(\log^{k}n)$, using bounded-fanin Boolean gates, unbounded-fanin $\mathsf{AND}$$\mathsf{OR}$ gates, and threshold gates, respectively.
We then characterize the exact computational power of latent thought in the parallel computation regime.
@@
For each $k\in\mathbb{N},$ it holds that
$\mathsf{Loop}[\log^{k}n,\,\mathsf{poly}(n),1\ \textup{(resp.\ }\log n)]$
$=\mathsf{CT}[\log^{k}n,\,\mathsf{poly}(n),1\ \textup{(resp.\ }\log n)]$
$=\mathsf{AC}^{k}\ (\textup{resp.\ }\mathsf{TC}^{k}).$
@@
The inclusion from circuits to latent thought follows from Theorem 3.10. For the converse inclusion, we build on the arguments of prior work (Merrill and Sabharwal, 2023; Li et al., 2024), which show that a fixed-depth Transformer block under finite precision is contained in $\mathsf{AC}^{0}$ (or $\mathsf{TC}^{0}$ under logarithmic precision). We extend their analysis to the looped setting, which can be unrolled into a $\mathsf{TC}^{k}$ circuit by composing a $\mathsf{TC}^{0}$ block for $\log^{k}n$ iterations. ∎
Moreover, we establish an upper bound on the power of CoT in the parallel computation regime. This limitation arises from the inherently sequential nature of CoT.
@@
For each $k\in\mathbb{N},$ it holds that
$\mathsf{CoT}[\log^{k}{n},\mathsf{poly}(n),\log{n}]\subseteq\mathsf{TC}^{k-1}.$
@@
The total $\log^{k}n$ steps can be divided into $\log^{k-1}n$ blocks, each consisting of $\log n$ steps. Since $\mathsf{CoT}[\log n,\mathsf{poly}(n),\log n]\subseteq\mathsf{TC}^{0}$ (Li et al., 2024), each block with the previous block’s outputs fed as inputs to the next block can be simulated in $\mathsf{TC}^{0}$; iterating this over $\log^{k-1}n$ layers yields a circuit in $\mathsf{TC}^{k-1}$. ∎
Figure 3: The separation between latent thought and CoT for decision problems, under polylogarithmic iterations.
These results lead to a separation in expressive power under standard complexity assumptions, as illustrated in Figure 3.
@@
For each $k\in\mathbb{N}$, if $\mathsf{TC}^{k-1}\subsetneq\mathsf{NC}^{k}$, then
$\mathsf{CoT}[\log^{k}n,\mathsf{poly}(n),\log{n}]$ $\subsetneq\mathsf{Loop}[\log^{k}n,\mathsf{poly}(n),1],$
$\mathsf{CoT}[\log^{k}n,\mathsf{poly}(n),\log{n}]$ $\subsetneq\mathsf{CT}[\log^{k}n,\mathsf{poly}(n),1].$
@@
For each $k\in\mathbb{N}$, if $\mathsf{TC}^{k-1}\subsetneq\mathsf{TC}^{k}$, then
$\mathsf{CoT}[\log^{k}n,\mathsf{poly}(n),\log n]$ $\subsetneq\mathsf{Loop}[\log^{k}n,\mathsf{poly}(n),\log n],$
$\mathsf{CoT}[\log^{k}n,\mathsf{poly}(n),\log n]$ $\subsetneq\mathsf{CT}[\log^{k}n,\mathsf{poly}(n),\log n].$
@@
The claims follow directly from Theorem 3.12 and Lemma 3.13. The established separations of the complexity classes, as summarized in Fig. 3, show that latent thought reasoning enables efficient parallel solutions more effectively than CoT, which is inherently sequential.
4 CoT Enables Approximate Counting
In the previous section, we showed that for decision problems, latent thought can yield more efficient solutions than CoT. This naturally raises the question of whether latent thought is universally more powerful than CoT. While prior work has shown that CoT can be simulated by looped Transformer models for deterministic decision problems under deterministic decoding (Saunshi et al., 2025), we found that this result does not directly extend to probabilistic settings with stochastic decoding. Accordingly, we shift our focus from efficiency in terms of the number of reasoning steps to expressive capability under polynomially many iterations.
4.1 Preliminaries
Approximate counting.
Formally, let $\Sigma$ be a finite alphabet and let $R\subseteq\Sigma^{*}\times\Sigma^{*}$ be a relation. For an input $x\in\Sigma^{*}$, define $R(x):=\{\,y\in\Sigma^{*}\mid(x,y)\in R\,\},$ and the counting problem is to determine $|R(x)|$. A wide class of natural relations admits a recursive structure, which allows solutions to be constructed from smaller subproblems.
Definition 4.1 (Informal: Self-reducibility (Schnorr, 1976)).
A relation $R$ is self-reducible if there exists a polynomial-time procedure that, given any input $x$ and prefix $y_{1:k}$ (with respect to a fixed output order), produces a sub-instance $\psi(x,y_{1:k})$ such that every solution $z$ of $\psi(x,y_{1:k})$ extends $y_{1:k}$ to a solution of $R(x)$ (and conversely), i.e., $R\bigl(\psi(x,y_{1:k})\bigr)=\{z\mid\ \mathrm{concat}(y_{1:k},z)\in R(x)\,\}.$
While exact counting is intractable, there exist efficient randomized approximation algorithms (Karp and Luby, 1983).
@@ 4.1 Preliminaries
An algorithm is called a fully polynomial-time randomized approximation scheme (FPRAS) for a function $f$ if, for any $\varepsilon>0$ and $\delta>0$, it outputs an estimate $\hat{f}(x)$ such that
@@
The class of counting problems that admit an FPRAS is denoted by $\mathsf{FPRAS}$. Although randomized algorithms provide only probabilistic guarantees, they are often both more efficient and simpler than their deterministic counterparts, denoted by $\mathsf{FPTAS}$ (Definition C.11). For example, counting the number of satisfying assignments of a DNF formula admits an FPRAS based on Monte Carlo methods (Karp et al., 1989), whereas no FPTAS is known for this problem. Moreover, probabilistic analysis enables us to capture algorithmic behavior on typical instances arising in real-world applications (Mitzenmacher and Upfal, 2017).
Probabilistic models of computation.
In contrast to the deterministic models considered in the previous section, we now study probabilistic models that define a conditional distribution over output strings $y=(y_{1},\ldots,y_{m})\in\Sigma^{*}$. We consider autoregressive next-token prediction of the form
$p(y\mid x)=\prod_{i=1}^{m}p(y_{i}\mid x,y_{<i}),$
where the model is allowed to perform additional reasoning steps before producing each output token $y_{i}$. This formulation was first used to formalize CoT for language modeling by Nowak et al. (2024). For latent thought, reasoning iterations are performed entirely in hidden space; no linguistic tokens are sampled except for the output token $y_{i}$, as illustrated in Fig.  4. This definition is consistent with practical implementations (Csordás et al., 2024; Bae et al., 2025). Within this framework, we define complexity classes of probabilistic models, denoted by $\mathsf{pCoT}$, $\mathsf{pCT}$, and $\mathsf{pLoop}$, respectively. Formal definitions are in Section C.1.
Figure 4: Probabilistic models of computation with latent thought. Each output token $y_{i}$ is generated autoregressively, with computation permitted in a continuous latent space prior to each token.
4.2 Separation in Approximate Counting
We first analyze the expressivity of the token-level conditional prediction at each step, $p(y_{i}\mid x,y_{<i})$, and show that CoT is strictly more expressive than latent thought in this setting. The key distinction is whether intermediate computation permits sampling. CoT explicitly samples intermediate reasoning tokens, inducing stochastic computation and enabling the emulation of randomized algorithms. In contrast, latent thought performs only deterministic transformations in latent space, resulting in deterministic computation.
Lemma 4.3 (Informal).
Assume that $\mathsf{FPTAS}\subsetneq\mathsf{FPRAS}$ for self-reducible relations. There exists a self-reducible relation $R$ and an associated function $f:\Sigma^{*}\times\Sigma^{*}\to\mathbb{N}$ defined by $f(x,y_{<i})\coloneqq|\{\,z\in\Sigma^{*}:(x,y_{<i}z)\in R\,\}|$ such that CoT with polynomially many steps admits an FPRAS for $f$. Whereas, no latent thought with polynomially many iterations admits the same approximation guarantee.
Proof sketch.
For self-reducible relations, approximating the counting function $f$ on subproblems is polynomial-time inter-reducible with approximating $|R(x)|$ (Jerrum et al., 1986). If latent thought with polynomially many iterations admitted an FPTAS for $f$, then it would induce a deterministic FPTAS for $|R(x)|$, contradicting the assumption.∎